7. A 2.0-cm-tall object is 15 cm in front of a converging lens that has a 20 cm focal length. Use the thin lens equation to calculate the imagine position and height.
Answer #1

Given that,

Object height, ho = 2 cm

object distance, u = 15 cm

focal length, f = 20 cm

From thin lens equation,

1 / f = 1 / v + 1 / u

1 / v = 1 / 20 - 1 / 15

v = -60 cm

Image position is -60 cm.

Magnification m = -v / u = hi / ho

60 / 15 = hi / 2

hi = 8 cm

Height of image is 8 cm.